OpenStudy (anonymous):

Choose the equation of the line passing through the point (2, 6) and parallel to y = -3x - 4. y = 3x - 12 y = -3x y = -3x + 12 y = 3x

OpenStudy (anonymous):

(y-y1)=m(x-x1) parallel(has same slope) (y-6)=-3(x-2) y=-3x+12

OpenStudy (anonymous):

asaad is right i think

OpenStudy (anonymous):

so you just plug in the points and subtact?

OpenStudy (anonymous):

*subtract